UT 2020 Sep 29 - SOUL-SX/LUCI-1 Commissioning

OSA: J. Williiams

ISp: C. Bolyard

Mtn Manager: J. Prothro

ISA: B. Rothberg (undisclosed secret location), R. Rothberg (supervising from undisclosed location)
AO: D, Miller, J. Power (Tucson)


Night Plan:

  • OSCO
  • Cygnus A On-Axis
  • Cygnus A Off-Axis
  • HR 8799 ADI Tests
  • NGC 628 Region 2
  • Bin 4 (R > 15 AO Ref Star)

Problems:

  1. DIMM still not working
  2. ADI issue at high transit elevation
  3. Instrument config/LUCI timeout after 1060 seconds

Night Log:

UT 00:21 -starting LUCI GUIs, init all for both LUCIs (just in case)

L1 0001-0002 - 2.51sec dark

L2 0001-0002 - 2.51sec dark

UT 01:15 - enclosure open

UT 01:30 - pointing and collimation check

OSCO

UT 01:34 - sending OSCO preset to HD 229143

UT 01:41 - WFS arbitrator offline - inserted pause after L1 0008

-need to restart all the preocesses L1 0003-0008 DO NOT USE

UT 01:46 - resending preset for OSCO HD 229143

UT 01:55 - skip frames (L1 0018,0019)

Gopt is all over the place. Paused LUCI data collection

D. Miller turned off Gopt, now giving reasonable values

-reapplying Gopt, values oscillating again.

UT 01:58 - D. Miller putting gains bacl to working values, leaving Gopt off. No NCPA will be applied

UT 02:00 -restarting data collection, L1 0019 and earlier DO NOT USE

L0024 - moved manually back to first offset -5.0, +7.0 Absolute

D. Miller putting FLAO into intervention mode to skip the Gopt step. Will need to click down the list.

UT 02:07 - L1 0020-0042

-testing if can replicate the software failure earlier before proceeding to next target.

UT 02;10 - Gopt appears to be unusable tonight

Cygnus-A

UT 02:11 - sending preset to Cygnus A - on-axis

-Preset sent the AO Ref value as R=11 (which is what USNO labeled the core) - can't see anything in WFS. Binned 16x16, silver mirror in place. All light should be going into CCD47, looking as deep as possible.

-object is centered, bayside stages are centered

UT 02:21 - re-sending preset V2 (AO Ref shows R=17.5)

-ddn't see it, will try off-axis version

UT 02:25 - sending off-axis version of script R=11.6

Seeing 1" on guider/WFS

-AO Ref seen in CCD47, AO Ref Rmag reported as 13.7 (catalog listed as 11.6) - running in Bin 2

Loop Parameters:

estimated mag 13.7 Bin2, freq 856. Average pupil intensity ~26, total incoming light 5.4 photons/subapp

-gap in time, L1 0044 was after a pause which I removed. Had to manually hit GO

L1 0043-0068

UT 02:49 after -200",+122.0" offset failed to find the star (L1 0048 was last off)

-sending 0,0 relative offset using LUCI Telsvc, this kicked it back to the proper position for AO bayside stage TROUBLESHOOTING FIX FOR THIS SITUATION

-resuming LUCI script

Off field (L1 0060) we were guiding - so active optics running, measured FWHMs raning from 0.3-0.85"

L1 0062 - field star Bin 2 15" away from off-axis AO Ref star FWHM = 0.06" - can see nice airy rings

UT 03:32 - finished script

-sending 0,0 absolute offset

-sending relative -29.2, +12.2 offset - AO Ref didn't come back into CCD47 - bayside stage didn't move, timed out

-will try J. Power trick to recover (0,0 relative offset via telsvc)

-bayside stages went to correct position but no star

-moving to 0,0 DETXY absolute - star appeared

moving -28.4,-11.2 RA offset (relative) - THIS worked!

Something Failed with DETXY calculation

UT 03:45 - absorbing offset

-did test exposure with GEIRS - worked (peak pixel at center is 53k)

-loaded a new script (don't send preset) to do dither and change the Itimes

L1 0069-0073 - PSF of star used for AO Ref

-testing Gopt again this time with Bin2 - failed again

UT 03:55 - loading pair PSF similar separation (29

doing pointing check = all ok

WFS/Guider raeading 0.6"/0.55"

UT 04:02 - sending matched psf companion - WFS

AO Ref too faint

UT 04:07 - flipped above script, AO Ref star becomes target, target becomes AO Ref

Guider 0.56" / WFS shows 0.56"

UT 04:17 - sending 4th attmept to match AO Ref with what we had

Guider/WFS report 0.50"/0.63"

WFS reports Rmag = 13.5

L1 0074-0076 - stopped here - 90sec integrated - elongation may be due to flexure

L1 0077-0081 - 60 sec - 23.4" away vs 30.6" away

showing elongation - going to rotate by 15 degrees to see if we avoid vibration

UT 04:35 - sending dtheta + 15

-sending another 15 degree offset (now at +30)

-shell RIP'd, but AO thinks loops ara closed

-need to do AdSec stop, AdSec start

UT 04:45 -seems to be fixed

-made V6 of script with PA=30

UT 04:51 - sending preset

L1 0082 - still elongated - manually adjusting higher orders

UT 05:04 - adjusting +30 rotation in telsvc

UT 05:06 - adjusting -60, (above made it WAY worse)

-adding another -30

Core is round, changing the frequency (opening loops, setting frequency, closing loops)

L1 0083-0087

UT 05:28 - completed

HR8799 - ADI Test

UT 05:32 - sending collimation script

-guider/WFS 0.61"/ 0.58"

-insert pause after 1st and 2nd exposure

UT 05:38 - sending script - had AO Mag 30, fixed it with V2

Landed at 1103, 726

moving -1.185", +4.47" - had it set for RA/Dec

Moved it 2x5" in Det Y

at 1049,1091

Absorbing offsets

Subframe setup: -256x256 Xo=921, Yo=963

UT 05:54 - proceeding to 2nd stage - verifying with H2

-max counts at 51k - just barely understauration

UT 05:56 - starting

first offset -0.91" -0.66'

UT 05:58 - changing this to 1" box random dither from 2" - will load new script and skip to where we left

L1 0091-0095

pausing at L1 0095 - dithers seem to be off to the left

using telsvc to move 0,0 DetXY absolute

-clearly not zero

963,1112 - this is where it thinks its 0,0

1049,1091 SHOULD be the center

shift +1.29" , -0.315" relative DetXY

landed at 1040,1099

clicked absorb

Testing with 0,0 Absolute DetXY

now at 1032,1105

Could it be an absolute X Y is not working in DETXY?

sending RA Dec Absolute offset - now it vanished from the field

Sending DETXY 0,0 absolute offset

UT 06:14 - removed subframe and looks like target is now at 1193,1368

-sending 0,0 absolute DETXY offset

1191.1372

Now Subframe is: subwin SW 1 1063 1244 256 256

UT 06:21 - sent 0,0 DETXY absolute offsset, stayed put

-sneding 0.1,0.1 DETXY absolute

-sending 0,0 DETXY absolute now at 1184,1371

-sending -0.1", -0.1" DETXY absolute

Bayside stage in X didn't move, confirmed from image, it shifted in Y only

UT 06:26 - sending 0,0 DETXY absolute

-undid the Y move

-sending 0.2" X only ABSOLUTE DETXY

UT 06:31 - Starting from top of script

-sending 0,0 DETXY Absolute DETXY - now back to 1193,1372 (within 2 pixels of X)

-all these crazy jitters happened during transit (79 degrees)

-This type of program worked for S. Ertel, but that program at a lower elevation - it transited at 63 degree elevation

Should revisit this program again tomorrow night to verify findings, if it wanders off the frame continue with script to see if it wanders back

L1 0097-0118 - small offset size (within +/-0.5")

UT 06:52 - loading script with larger offsets

L1 0119 -0128

L0129-0137 back to small offsets

Info from AO system:

estimated mag 7.4 Bin1, freq 1700. Average pupil intensity ~140 total incoming light 170.2 photons/subapp

UT 07:11 Completed H2 filter

NGC 628 - Region 1

UT 07:16 - pointing check

UT 07:24 - collimation check

UT 07:24 - 0.7" on guider/WFS during colliimation check

UT 07:26 - guider/WFS in NGC 628 field 0.71"/0.66"

-WFS estimated mag 14.3 (script says R=13.4) Bin2, freq 765 Average pupil intensity ~17 total incoming light 3.4 photons/subapp
UT 07:33 -attempting to run Gopt
L1 0148 - small star next to saturataed WFS is 0.07" FWHM

L0149 - small star next to staurated WFS is 0.071" FWHM

UT 08:29 - guider 1.1" / WFS 0.9" (guiding during an offset)

UT 08:38 - guider 0.95" / WFS 0.89" (guiding during an offset)

Note: Gopt appears to be running OK now.

UT 08:47 - guider/WFS 0.8" (guiding on off field)

UT 08:56 - guider/WFS 1.07"/ 0.76 (guiding on off field)

UT 09:00 - guider/WFS 0.86"/0.84" (guiding in off field)
L1 0138-0170

UT 09:05 - completed NGC 628 Region 1

Performance Testing

UT 09:10 - moving to Star #1 02:59:58 +19:56:40 R=15.3

Star 1 R=15.3, WFS Rmag = 17.2 ; 600Hz - safe skipped

-2nd attempt safe skipped

-thrid attempt to close the loop

estimated mag 17.2 Bin4 freq 613 Average pupil intensity ~9 total incoming light 1.5 photons/subapp
-attempting a 4th time to cose the loop

Insturment config timed out after 1600 seconds - config is in place and preset is still valid

WFS showing 1.1" guider, 0.9" WFS

- gave up

UT 09:31 - sending preset to Star #2 02:59:54 +20:04:23 R=15.47

-guider reporting seeing ~ 2"

WFSmag reporting R=17.4

UT 09:38 Holding in place - wating for conditions to improve....

UT 09:44 - seeing variable

UT 09:53 - canceled AO preset, fell back to Active preset, monitoring Guider

UT 09:58 - resending preset to star....

seeing now about 0.85" - J. Power closing the loop

WFS reporting R=17.0

-safe skip

UT 10:10 - still trying to close the loop

2.51 sec gets about 76 counts

-preset still shows green even though loop is closed

-after injecting the backdoor script the preset finished...very odd!

UT 10:15 - starting data collection

AO info:

estimated mag 16.9-17.0 Bin4 freq 650 Average pupil intensity ~8 total incoming light 1.5 photons/subapp

UT 10:25 - safe skipped but loops did not open

-GS wandered off the field

Cancelling preset

UT 10:26 - resending preset - aborted exposure -still can't cancel AO preset

UT 10:28 - resending preset

UT 10:39 - starting data collection

estimated mag 16.9-17.0 Bin4 freq 640 Average pupil intensity ~8 total incoming light 1.5 photons/subapp

-safe skipped, had to abort

UT 10:41 - resendng the preset to check the seeing

-not a triangular psf like bfore....

Seeing from guider/WFS 0.9"/0.8"

UT 10:46 - updating AOS table - mistake was found - set too high a frequency for Bin4 R~17

UT 10:49 - canceling AO preset

UT 10:49 - sending preset

changed AO mag in script to R=17.0 instead of 15.2

UT 10:54 - sending preset

-guider 1.2" WFS 0.9"

estimated mag 17.3 Bin4 freq 320 Average pupil intensity ~14 total incoming light 2.4 photons/subapp, Failed on close loop

UT 11:08 - attempting to close loop again

Additional values updated, should improve optimize gain

UT 11:14 - taking 10secondtest exposure ~1100 counts, looks elongated

UT 11:16 - collecting data

estimated mag 17.6 Bin4 freq 321 Average pupil intensity ~14 total incoming light 1.8 photons/subapp
-new config - gains are low, almost no corrections appliied

UT 11:28 - taking 30" offset in X in DET XY absolute for dark

UT 11:40 - moving to Star 4 R=15.5 03:20:22 +20:34:31

-re-rotator error

UT 11:43 - flipped field 180 degrees - sending revised script

WFS/Guider show 0.8" seeing

WFS reports R=16.6 (script/catalog Show R=15.05)

Rotator = 130 PA need to change PA, manually using telvc to -30 deg from current position, Riotator is at 100.

UT 11:54 - moving another -30 degrees.

Rotator is at 70 degrees

Moving +30 degrees, rotator = 160 degrees

Now sending -180 absolute, rotator is now 310

UT 12:03 - fixing fieldstop it shiftd quite a bit

UT 12:08 - finally aligned fieldstop - resending original preset PA=0

Seeing 0.8" on guider/WFS

WFS reports R=16.4, vs R=15.5 from catalog

AO info:

Rerotator error so flipped 180, but then railed camera lens, flipped back and rerotator error gone. (bin4 target)

estimated mag 16.4 Bin3 freq 260 Average pupil intensity ~17 total incoming light 3.4 photons/subapp,

No obvious minimum in optimize gains plots first run through so tweaking the ranges and rerunning

UT 12:26 - LUCI insturment config timeout

UT 12:27 - backround getting, well into nautical twilight

UT 12:28 - stopping on-sky, will do a few cals
Calibrations

UT 12:48 - Ks flat for N30

L1 0183-0187 lights off

L1 0188-0192 lights on

UT 12:55 - finished, putting LUCIs in safe config

-- BarryRothberg - 28 Sep 2020
Topic revision: r2 - 29 Sep 2020, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback